Find Continuity of
f(x,y)=y/√(x^2+y^2). at (x,y)=(0,0)
1
Expert's answer
2021-07-19T05:51:50-0400

"\lim\nolimits_{{(x,y)\to(a,b)}} f(x,y)= f(a,b)"  is the definition of continuity. Now take the limit from the x-axis

"\lim\nolimits_{{(x,y)\to(0,0)}} f(x,0)=\lim\nolimits_{{(x,y)\to(0,0)}} \frac{0}{\sqrt{(x^2+0)}}=0\\"

Now take the limit from x=y

"\lim\nolimits_{{(x,y)\to(0,0)}} f(x,x)=\lim\nolimits_{{(x,y)\to(0,0)}} \frac{x}{\sqrt{(x^2+x^2)}}\\
\lim\nolimits_{{(x,y)\to(0,0)}} \frac{x}{\sqrt{(2x^2)}}= \frac{1}{\sqrt{2}}\\"

Limit from the x-axis is 0 but the limit from x=y line is "\frac{1}{\sqrt{2}}" therefore the limit does not exist which means at (0,0) the funtion f(x,y) is discontinuous since limit f(x,y) does not equal to f(0,0)